how to solve the equation x^2=3-2x by completing the square

Found 2 solutions by richwmiller, nerdybill:
Answer by richwmiller(17219)

x^2+2x=3
x^2+2x+1=4
(x+1)^2=4
x+1=2
x=1
x+1=-2
x=-3
also could have been factored
(x+3)*(x-1)=0

Answer by nerdybill(7384)

x^2=3-2x
x^2+2x = 3
take (1/2)b, square the result and add to both sides:
x^2+2x+1 = 3+1
(x+1)(x+1) = 4
(x+1)^2 = 4
x+1 = sqrt(4)
x+1 = (+-)2
x = (+-)2 - 1
x = {-3, 1}